# Why Transformers Need Adam: A Hessian Perspective

### v2

@@ Abstract
SGD performs worse than Adam by a significant margin on Transformers, but the reason remains unclear. In this work, we provide an explanation of SGD’s bad performance on Transformers through the lens of Hessian: (i) Transformers are “heterogeneous”: the Hessian spectrum across parameter blocks vary dramatically, a phenomenon we call “block heterogeneity"; (ii) Heterogeneity hampers SGD: SGD performs badly on problems with block heterogeneity. To validate that heterogeneity hampers SGD, we check various Transformers, CNNs, MLPs, and quadratic problems, and find that SGD works well on problems without block heterogeneity but performs badly when the heterogeneity exists. Our initial theoretical analysis indicates that SGD performs poorly because it applies one single learning rate to all blocks, which cannot handle the heterogeneity among blocks. This limitation could be ameliorated if we use coordinate-wise learning rates, as designed in Adam. [^1] Our code is available at https://github.com/zyushun/hessian-spectrum.
@@ 1 Introduction
Transformers [84] have become a major workhorse behind much remarkable progress in AI development (e.g., [1]). However, the understanding of Transformer training remains limited. For instance, Transformer training largely relies on the Adam optimizer [45, 56]. In contrast, stochastic gradient descent with momentum (SGD) [^2] We introduce the update rules of Adam(W) and SGD in Appendix C.1., the de-facto optimizer for convolution neural networks (CNNs) [48], performs poorly on Transformers (e.g., Figure LABEL:fig_blockwise_spectrum). Yet, the reasons behind this poor performance remain unclear. Understanding why SGD performs badly on Transformers is an intriguing question. First, from a theoretical perspective, this can help us better understand the training of Transformers and more generally, neural networks. Second, from a computational perspective, the understanding may inspire the design of better algorithms for training neural networks.
@@
In this work, we explore the cause for SGD’s deficiency through the lens of Hessian. We start by investigating the full Hessian spectrum of Transformers, i.e., the full eigenvalue density of Hessian (see Figure LABEL:fig_full_spectrum). By theory, the full Hessian spectrum largely determines the behavior of gradient-based methods [63, 33, 80, 37], so we suspect it may also help explain SGD’s bad performance on Transformers. Using tools from numerical linear algebra [7], we empirically compare the full spectra of CNNs (where SGD is on par with Adam) and those of Transformers (where SGD largely lags behind Adam). Unfortunately, as shown in Figure LABEL:fig_full_spectrum, the spectra for CNNs and Transformers are often largely similar despite the different optimizer behaviors. As such, we have not identified critical features in the full Hessian spectra associated with SGD’s bad performance on Transformers. To reveal the cause, a more fine-grained investigation into the Hessian is needed.
@@
What would cause SGD to perform badly on Transformers but not on CNNs? By dissecting the structures of CNNs and Transformers, we notice that CNNs are constructed by the repetitive stacking of similar parameter blocks (convolution layers), while Transformers involve the non-sequential stacking of disparate parameter blocks (e.g. Query, Key, Value, Output projection blocks in attention and MLP layers). We hypothesize that these architectural differences might lead to different optimization properties. Intuitively, disparate parameter blocks contribute differently to the overall loss. So each block might benefit from a specialized treatment by optimizers, a flexibility offered by Adam but not by SGD. This observation motivates us to investigate the Hessian spectrum of each parameter block, which we refer to as the blockwise Hessian spectrum.
@@
By inspecting the blockwise Hessian spectrum, we discover a possible explanation for SGD’s bad performance: the “heterogeneity” inherent in Transformers. We provide both empirical and theoretical evidence to support this explanation. Our contributions can be summarized as follows:
@@
  • Failure mode of SGD on Transformers. We explain why SGD underperforms Adam on Transformers by examining the blockwise Hessian spectrum. First, we identify a phenomenon called ’block heterogeneity’, which refers to the large differences in the Hessian spectra across parameter blocks. This block heterogeneity is observed in all examined Transformers but not in CNNs. Second, we verify that block heterogeneity hinders SGD. Across various Transformers, CNNs, and MLPs, we show that SGD consistently performs poorly on problems with block heterogeneity but works well otherwise.
@@
  • Theoretical results on quadratic models. We construct convex quadratic problems with and without block heterogeneity and find that gradient descent (GD) performs significantly worse than Adam on problems with block heterogeneity but comparably otherwise. Our theoretical analysis shows that GD struggles on quadratic problems with block heterogeneity, suggesting that using a single learning rate for all blocks is a key reason for GD’s bad performance. The deficiency can be mitigated by assigning different learning rates across blocks, as Adam does.
@@
We emphasize that we do not claim block heterogeneity is the only cause for SGD’s bad performance, but just that it is at least one important cause. We verify, both empirically and theoretically, that SGD performs poorly when block heterogeneity is present.
@@ 2.2 Full Hessian Spectrum Is Not Informative Enough
We study the full Hessian spectrum of Transformers for two reasons. First, as stated in Section 1, the Hessian spectrum significantly influences the behavior of gradient methods [63]. Second, previous research shows that the Hessian spectrum provides insights into neural network phenomena, like BatchNorm’s effect on training speed [32]. Therefore, we hypothesize that the Hessian spectrum may also explain SGD’s bad performance on Transformers.
We compare the full Hessian spectra of CNNs (where SGD works well) and those of Transformers (where SGD struggles), as shown in Figure LABEL:fig_full_spectrum. Unfortunately, the results suggest that the full Hessian spectrum alone may not suffice to explain why SGD struggles on Transformers. We elaborate as follows. The primary information in the spectrum lies in its (A) dispersion, (B) shape, and (C) evolution during training. Regarding (A), we observe that the eigenvalues are dispersed similarly across different models, with no notably large outlier for Transformers. Thus, dispersion does not seem to be related to the bad performance of SGD. We further investigate (B) and (C). For all CNNs and Transformers in Figure LABEL:fig_full_spectrum, we observe similar phenomena: the spectrum’s shape is approximately symmetrical around 0 at initialization. As training proceeds, the majority of negative eigenvalues disappear, and the shape evolves into a combination of a “bulk” and some “outliers”. Since the spectral shape and evolution are quite similar for both Transformers and CNNs, they cannot explain the incompetence of SGD on Transformers either. In summary, we have not identified any critical phenomena in the full Hessian spectra that can be linked to SGD’s bad performance on Transformers.
@@ 2.3 Main Findings Through Blockwise Hessian Spectra
What other factors could cause SGD to struggle on Transformers but not on CNNs? We identify some critical features that have been overlooked in the full Hessian spectrum analysis above.
@@
The results in Figure LABEL:fig_blockwise_heatmap showed a new phenomenon: for all Transformers we checked, the blockwise Hessian spectra are largely different across blocks. In the following, we refer to this phenomenon as “block heterogeneity". In contrast, the blockwise Hessian spectra of CNNs are similar and the block heterogeneity is not observed. We refer to this phenomenon as “block homogeneity". These results indicate that block heterogeneity is informative in distinguishing CNNs and Transformers. Intuitively, the block-homogeneity in CNNs comes from the repetitively similar convolution layers, while the block-heterogeneity in Transformers stems from the non-sequential stacking disparate layers such as Query, Value, and MLPs. In the following, we will show that the block heterogeneity is strongly correlated with SGD’s bad performance on Transformers.
2.4 SGD Performs Badly on Various Tasks with Block Heterogeneity
Figure LABEL:fig_blockwise_spectrum and LABEL:fig_blockwise_heatmap have shown that (1) SGD performs badly on Transformers. (2) Transformers have block-heterogeneity. Now we further link block heterogeneity to SGD’s bad performance on non-Transformer models. This would directly establish a connection between block heterogeneity and SGD’s bad performance, without going through Transformers or attention blocks as an intermediary. We consider one man-made example and one real-world example.
Example 1: A man-made MLP. We consider a 4-layer MLP on MNIST and change the degree of heterogeneity by scaling each layer by constant $c$. Figure LABEL:fig_mlp_gap (a) shows SGD fails as heterogeneity grows.
@@ 2.6 Implication on Choosing SGD or Adam
The above findings show that SGD can perform badly on various architectures. This leads to an intriguing question: Can we predict the bad performance of SGD before training begins?
@@ 4 Conclusion
In this work, we explore why SGD performs badly on Transformers. By numerically exploring various neural nets and quadratic problems, we establish a phenomenon called block heterogeneity in Hessian and link it to the deficiency of SGD. Initial theory is provided to support the claim.
@@ Appendix A Related Works
On the deficiency of SGD on Transformers
There is an active line of works that explores why SGD performs badly on Transformers. One representative hypothesis is that SGD cannot handle the heavy-tailed stochastic noise in language tasks [106]. However, Chen et al. [14], Kunstner et al. [46] reported that the gap between Adam and SGD maintains even in the full-batch case with no stochasticity, so there might be other reasons. Further, SGD performs poorly on Vision Transformers on ImageNet (See Figure LABEL:fig:nlp_figure. Also see [93] for more evidence), so the data modality (e.g., language or vision tasks) might not be as crucial as the architecture. [105] showed that NLP tasks have “unbounded smoothness" issue and SGD with gradient clipping performs better than SGD in this case. Although clipping is an effective trick, we still observe a huge gap between clipped SGD and Adam  [^4] For all NLP tasks, clipping is performed immediately after backpropagation. So in Figure LABEL:fig:nlp_figure, SGD in NLP tasks essentially refers to clipped SGD. . So there might be other reasons that hamper SGD. Different from these works, we find SGD performs badly because it uses one single learning rate for all blocks, which cannot handle the Hessian heterogeneity among blocks.
@@
There was once a long-standing debate on the possible divergence of Adam [72]. The convergence for the unmodified versions is later established in [78, 108] for RMSprop and Adam. More convergence analyses of general adaptive gradient methods are listed later in this section. We here focus on the literature that explores the benefit of Adam. Xie et al. [94] show that Adam can help avoid saddle points, which is an orthogonal direction to this work. Wang et al. [85], Crawshaw et al. [19], Li et al. [50] show that Adam and its variant outperform SGD under relaxed smoothness conditions, based on the intuition that Adam can adaptively change its learning rate along iteration (over time). We pointed out that the theory is not complete: even for quadratic functions where the smoothness is fixed, SGD sometimes performs badly while Adam works well (Figure LABEL:fig_quadratic). This indicates that the benefit of Adam is not merely due to its ability to adaptively change the learning rate (over time), and there are other reasons for Adam’s success. We show that an important benefit of Adam is its ability to handle the heterogeneity across blocks (over space).
@@
Several solutions have been proposed. Liu et al. [55] numerically observed that adaptive gradient methods can (partly) overcome gradient vanishing by giving “consistent update magnitude", while it seems unclear how consistent update magnitude would help optimization in principle. Researchers further develop training tricks such as warmup learning rate [54, 95], temperature scaling [65], better initialization [103, 41, 88, 5, 97], and variants of Layer Norm [64, 89, 95, 88, 24]. Recent researchers also suggest using z-loss regularization [17, 96] and tuning hyperparameters of Adam [108, 91]. All these tricks can help mitigate gradient explosion or vanishing. Nevertheless, training large-scale Transformers remains challenging [107, 101, 91, 62, 17]. Different from all aforementioned works, we investigate the training difficulties of Transformers through the eigenvalues of Hessian. We establish a strong correlation between the blockwise Hessian spectra of Transformers and the poor performance of SGD. We realize that our attempt is just a first step towards understanding Transformer training, and we believe there is rich information hidden in Hessian and we leave more fine-grained analysis as future works.
@@ Appendix B More Results and Discussions
Here, we show the performance comparison of AdamW and SGD on different models. All the vision models are trained on ImageNet. Language models are trained on different English corpus. See Appendix D.1 for more implementation details.

### v3

@@ Abstract
SGD performs worse than Adam by a significant margin on Transformers, but the reason remains unclear. In this work, we provide an explanation through the lens of Hessian: (i) Transformers are “heterogeneous”: the Hessian spectrum across parameter blocks vary dramatically, a phenomenon we call “block heterogeneity"; (ii) Heterogeneity hampers SGD: SGD performs worse than Adam on problems with block heterogeneity. To validate (i) and (ii), we check various Transformers, CNNs, MLPs, and quadratic problems, and find that SGD can perform on par with Adam on problems without block heterogeneity, but performs worse than Adam when the heterogeneity exists. Our initial theoretical analysis indicates that SGD performs worse because it applies one single learning rate to all blocks, which cannot handle the heterogeneity among blocks. This limitation could be ameliorated if we use coordinate-wise learning rates, as designed in Adam. [^1] Our code is available at https://github.com/zyushun/hessian-spectrum.
@@ 1 Introduction
Transformers [84] have become a major workhorse behind AI development (e.g., [1]). However, the understanding of Transformer training remains limited. For instance, Transformer training largely relies on the Adam optimizer [45, 56]. In contrast, stochastic gradient descent with momentum (SGD) [^2] We introduce the update rules of Adam(W) and SGD in Appendix C.1., the de-facto optimizer for convolution neural networks (CNNs) [48], performs significantly worse than Adam on Transformers (e.g., Figure LABEL:fig_blockwise_spectrum). Yet, the reasons behind this performance gap remain unclear. Understanding why SGD performs worse than Adam on Transformers is an intriguing question. First, from a theoretical perspective, this can help us better understand the training of Transformers and more generally, neural networks. Second, from a computational perspective, the understanding may inspire the design of better algorithms for training neural networks.
@@
In this work, we explore why SGD largely underperforms Adam on Transformers through the lens of Hessian. We start by investigating the full Hessian spectrum of Transformers, i.e., the full eigenvalue density of Hessian (see Figure LABEL:fig_full_spectrum). By theory, the full Hessian spectrum largely determines the behavior of gradient-based methods [63, 33, 80, 37], so we suspect it may also help explain SGD’s unsatisfactory performance. Using tools from numerical linear algebra [7], we empirically compare the full spectra of CNNs (where SGD is on par with Adam) and those of Transformers (where SGD largely lags behind Adam). Unfortunately, as shown in Figure LABEL:fig_full_spectrum, the spectra for CNNs and Transformers are often largely similar despite the different optimizer behaviors. As such, we have not identified critical features in the full Hessian spectra associated with the gap between Adam and SGD on Transformers. To reveal the cause, a more fine-grained investigation into the Hessian is needed.
@@
What would cause SGD to perform significantly worse than Adam on Transformers, but not on CNNs? By dissecting the structures of CNNs and Transformers, we notice that CNNs are constructed by the repetitive stacking of similar parameter blocks (convolution layers), while Transformers involve the non-sequential stacking of disparate parameter blocks (e.g. Query, Key, Value, Output projection blocks in attention and MLP layers). We hypothesize that these architectural differences might lead to different optimization properties. Intuitively, disparate parameter blocks contribute differently to the overall loss. So each block might benefit from a specialized treatment by optimizers, a flexibility offered by Adam but not by SGD. This observation motivates us to investigate the Hessian spectrum of each parameter block, which we refer to as the blockwise Hessian spectrum.
@@
By inspecting the blockwise Hessian spectrum, we discover a possible explanation for why SGD is worse: the “heterogeneity” inherent in Transformers. We provide both empirical and theoretical evidence to support this explanation. Our contributions can be summarized as follows:
@@
  • Why SGD underperforms Adam on Transformers. We explain why SGD is worse than Adam on Transformers by examining the blockwise Hessian spectrum. First, we identify a phenomenon called “block heterogeneity", which refers to the large differences in the Hessian spectra across parameter blocks. This block heterogeneity is observed in all examined Transformers but not in CNNs. Second, we verify that block heterogeneity hinders SGD. Across various Transformers, CNNs, and MLPs, we show that SGD consistently performs worse than Adam on problems with block heterogeneity but can perform similarly to Adam, otherwise.
@@
  • Theoretical results on quadratic models. We construct convex quadratic problems with and without block heterogeneity and find that gradient descent (GD) largely underperforms Adam on problems with block heterogeneity, but can perform comparably otherwise. Our theoretical analysis shows that GD can be slower than Adam on quadratic problems with block heterogeneity. We point out GD is slower than Adam because it uses a single learning rate for all blocks. The deficiency can be mitigated by assigning different learning rates across blocks, as Adam does.
@@
We emphasize that we do not claim block heterogeneity is the only cause for the performance gap between Adam and SGD, but just that it is at least one important cause. We verify, both empirically and theoretically, that SGD underperforms Adam when block heterogeneity is present.
@@ 2.2 Full Hessian Spectrum Is Not Informative Enough
We study the full Hessian spectrum of Transformers for two reasons. First, as stated in Section 1, the Hessian spectrum significantly influences the behavior of gradient methods [63]. Second, previous research shows that the Hessian spectrum provides insights into neural network phenomena, like BatchNorm’s effect on training speed [32]. Therefore, we hypothesize that the Hessian spectrum may also explain why SGD largely lags behind Adam on Transformers.
We compare the full Hessian spectra of CNNs (where SGD performs similarly to Adam) and those of Transformers (where SGD underperforms Adam), as shown in Figure LABEL:fig_full_spectrum. Unfortunately, the results suggest that the full Hessian spectrum alone may not suffice to explain the gap between Adam and SGD on Transformers. We elaborate as follows. The primary information in the spectrum lies in its (A) dispersion, (B) shape, and (C) evolution during training. Regarding (A), we observe that the eigenvalues are dispersed similarly across different models, with no notably large outlier for Transformers. Thus, dispersion does not seem to be related to why SGD is worse than Adam. We further investigate (B) and (C). For all CNNs and Transformers in Figure LABEL:fig_full_spectrum, we observe similar phenomena: the spectrum’s shape is approximately symmetrical around 0 at initialization. As training proceeds, the majority of negative eigenvalues disappear, and the shape evolves into a combination of a “bulk” and some “outliers”. Since the spectral shape and evolution are quite similar for both Transformers and CNNs, they cannot explain why SGD is worse than Adam on Transformers, either. In summary, we have not identified any critical phenomena in the full Hessian spectra that can be linked to the performance gap between Adam and SGD on Transformers.
@@ 2.3 Main Findings Through Blockwise Hessian Spectra
What other factors could cause SGD to perform significantly worse than Adam on Transformers but not on CNNs? We identify some critical features that have been overlooked in the full Hessian spectrum analysis above.
@@
The results in Figure LABEL:fig_blockwise_heatmap showed a new phenomenon: for all Transformers we checked, the blockwise Hessian spectra are largely different across blocks. In the following, we refer to this phenomenon as “block heterogeneity". In contrast, the blockwise Hessian spectra of CNNs are similar and the block heterogeneity is not observed. We refer to this phenomenon as “block homogeneity". These results indicate that block heterogeneity is informative in distinguishing CNNs and Transformers. Intuitively, the block-homogeneity in CNNs comes from the repetitively similar convolution layers, while the block-heterogeneity in Transformers stems from the non-sequential stacking disparate layers such as Query, Value, and MLPs. In the following, we will show that the block heterogeneity is strongly correlated with the performance gap between SGD and Adam on Transformers.
2.4 SGD Performs Worse than Adam on Various Tasks with Block Heterogeneity
Figure LABEL:fig_blockwise_spectrum and LABEL:fig_blockwise_heatmap have shown that (1) SGD is worse than Adam on Transformers. (2) Transformers have block-heterogeneity. Now we further link block heterogeneity to SGD’s unsatisfactory performance on non-Transformer models. This would directly establish a connection between “block heterogeneity" and “why SGD is worse than Adam", without going through Transformers or attention blocks as an intermediary. We consider one man-made example and one real-world example.
Example 1: A man-made MLP. We consider a 4-layer MLP on MNIST and change the degree of heterogeneity by scaling each layer by constant $c$. Figure LABEL:fig_mlp_gap (a) shows SGD gradually performs worse than Adam as heterogeneity grows.
@@ 2.6 Implication on Choosing SGD or Adam
We have shown that SGD can largely underperform Adam on various architectures. This leads to an intriguing question: Can we predict the incompetence of SGD before the training begins?
@@ 4 Conclusion
In this work, we explore why SGD largely underperforms Adam on Transformers. we establish a phenomenon called block heterogeneity in Hessian and link it to the performance gap between Adam and SGD. Initial theory is provided to support the claim.
@@ Appendix A Related Works
On the unsatisfactory performance of SGD on Transformers
There is an active line of works that explores why SGD performs significantly worse than Adam on Transformers. One representative hypothesis is that SGD cannot handle the heavy-tailed stochastic noise in language tasks [106]. However, Chen et al. [14], Kunstner et al. [46] reported that the gap between Adam and SGD maintains even in the full-batch case with no stochasticity, so there might be other reasons. Further, SGD performs worse than Adam on Vision Transformers on ImageNet (See Figure LABEL:fig:nlp_figure. Also see [93] for more evidence), so the data modality (e.g., language or vision tasks) might not be as crucial as the architecture. [105] showed that NLP tasks have “unbounded smoothness" issue and SGD with gradient clipping performs better than SGD in this case. Although clipping is an effective trick, we still observe a huge gap between clipped SGD and Adam  [^4] For all NLP tasks, clipping is performed immediately after backpropagation. So in Figure LABEL:fig:nlp_figure, SGD in NLP tasks essentially refers to clipped SGD. , so there might be other reasons. Different from these works, we find SGD underperforms Adam because it uses one single learning rate for all blocks, which cannot handle the Hessian heterogeneity among blocks.
@@
There was once a long-standing debate on the possible divergence of Adam [72]. The convergence for the unmodified versions is later established in [78, 108] for RMSprop and Adam. More convergence analyses of general adaptive gradient methods are listed later in this section. We here focus on the literature that explores the benefit of Adam. Xie et al. [94] show that Adam can help avoid saddle points, which is an orthogonal direction to this work. Wang et al. [85], Crawshaw et al. [19], Li et al. [50] show that Adam and its variant outperform SGD under relaxed smoothness conditions, based on the intuition that Adam can adaptively change its learning rate along iteration (over time). We pointed out that the theory is not complete: even for quadratic functions where the smoothness is fixed, SGD sometimes performs largely worse than Adam (Figure LABEL:fig_quadratic). This indicates that the benefit of Adam is not merely due to its ability to adaptively change the learning rate (over time), and there are other reasons for Adam’s success. We show that an important benefit of Adam is its ability to handle the heterogeneity across blocks (over space).
@@
Several solutions have been proposed. Liu et al. [55] numerically observed that adaptive gradient methods can (partly) overcome gradient vanishing by giving “consistent update magnitude", while it seems unclear how consistent update magnitude would help optimization in principle. Researchers further develop training tricks such as warmup learning rate [54, 95], temperature scaling [65], better initialization [103, 41, 88, 5, 97], and variants of Layer Norm [64, 89, 95, 88, 24]. Recent researchers also suggest using z-loss regularization [17, 96] and tuning hyperparameters of Adam [108, 91]. All these tricks can help mitigate gradient explosion or vanishing. Nevertheless, training large-scale Transformers remains challenging [107, 101, 91, 62, 17]. Different from all aforementioned works, we investigate the training difficulties of Transformers through the eigenvalues of Hessian. We establish a strong correlation between “the blockwise Hessian spectra of Transformers" and “why SGD largely underperforms Adam on Transformers". We realize that our attempt is just a first step towards understanding Transformer training, and we believe there is rich information hidden in Hessian and we leave more fine-grained analysis as future works.
@@ Appendix B More Results and Discussions
Here, we show the performance comparison of AdamW and SGD on different models. All the vision models are trained on ImageNet. Language models are trained on different English corpus. We grid-search the learning rates for SGD and Adam under the same budget and report the best result for each optimizer. See Appendix D.1 for more implementation details.